1) what's the mass in 2.6 mol of C8H12 (Which is called windowpane) MM = 108.18g/ Mol
2) how many moles of magnesium is 3.01 x10 22 atoms of magnesium ? Na + 6.02x 10 23
1
Expert's answer
2020-12-18T08:37:05-0500

1.FindthemassofC8H12m=nMr=2.6∗108=280.8gMr(C8H12)=8Ar(C)+12Ar(H)=8∗12+12=1082FindtheamountofsubstanceofMg.n=N/Na=3.01∗102

2/6.02∗102

3=0.05mol